how do you write an quadratic equation that have given number7,-7 as the solution
Answer by CharlesG2(834)

how do you write an quadratic equation that have given number7,-7 as the solution
(x + 7)(x - 7) = 0 (as you can see solutions would be plus or minus 7)
x^2 - 7x + 7x - 49 = 0 (used foil)
x^2 - 49 = 0